OpenStudy (anonymous):

Could someone help I'll give you a medal please estimate 32,690- 8,044

OpenStudy (shubhamsrg):

Hint: answer is a 5 digit number

OpenStudy (shubhamsrg):

2nd digit = 4th digit 3rd digit = 5th digit

OpenStudy (shubhamsrg):

sum of digits =22

OpenStudy (shubhamsrg):

first 3 digits are all even and form an arithmetic progression

OpenStudy (anonymous):

I don't understand

OpenStudy (shubhamsrg):

Am quite sure you don't! :)

OpenStudy (anonymous):

32,690- 8,044 = 32700 - 8040 = 24660

OpenStudy (anonymous):

This was one of my student question...

OpenStudy (shubhamsrg):

And, you've done the calculation wrong.

OpenStudy (anonymous):

"estimate"

OpenStudy (shubhamsrg):

I see. So you;re approximating..hmm

OpenStudy (anonymous):

Yes not asking for exact

OpenStudy (anonymous):

Thanks@SgReborn

OpenStudy (anonymous):

But i still do not understand

OpenStudy (anonymous):

Pls could you tell me the method you used

OpenStudy (anonymous):

Approximation = Estimation by definition

OpenStudy (anonymous):

Please, what Is the solution and what method did you use

OpenStudy (anonymous):

Anything lower than 4 round off to 0, anything more than 5 round off to next place. In your case was, 32,690- 8,044, where the nearest estimate or approximation could take place in from tens to hundreds for 32,690 to 32,700, while 8,044 from ones to tens place. In addition since there is any large difference to round off by an increase of ten, as it was offset by 4 down.

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

Here is your medal